The total electric flux from a cubical box 33.0 cm
Answer #1

Here ,

side of cube , a = 33 cm

a = 0.33 m

flux through each side = 1850 N.m^2/C

let the charge enclosde is q

Using Gauss law

Total flux = Qin/Epsilon

1850 N.m^2/C = Qin/(8.854 *10^-12)

solving for Qin

Qin = 1.64 *10^-87 C

Qin = 16.4 nC

the charge enclosed in the cube is 16.4 nC